OpenStudy (anonymous):

Solve for x in [0,2pi] Cos^2x-sin^2x=0

sam (.sam.):

Put it in tangent form then solve it using the range given

OpenStudy (anonymous):

How do I put it in tangent form?

sam (.sam.):

Like this, \[\cos^2(x)-\sin^2(x)=0\] \[\cos^2(x)=\sin^2(x)\] \[1=\frac{\sin^2(x)}{\cos^2(x)} \] \[ \tan^2(x)=1\] \[\tan(x)=\pm 1\]

sam (.sam.):

Now we have "\(\pm \)" so we'll use all the quadrants 

OpenStudy (tkhunny):

?? \(\cos^{2}(x) - \sin^{2}(x) = \cos(2x)\). Why do it some hard way?

sam (.sam.):

\[\tan(x)=1 \\ x=45 \] \[x=180-45=135 \\ x=180-45=225 \\ x=360-45=315\] 4 answers

OpenStudy (anonymous):

Awesome, thank you so much. @.Sam.

OpenStudy (tkhunny):

Unfortunately, the "tangent" method introduced Domain restrictions. If cos(x) = 0 had been a solution to the original equation, you would have discarded it with the division to tangent. Always keep the Domain in mind. Always worry about changing it.

OpenStudy (anonymous):

How would I do sin3x+sinx=0 with the same domain?

sam (.sam.):

sin(3x)+sin(x)=0?

OpenStudy (anonymous):

Yes

terenzreignz (terenzreignz):

Looks horrible... thankfully, we have our ever useful identities... unless you already know what sin(3x) is... Do you, by any chance? :D

OpenStudy (anonymous):

No

OpenStudy (anonymous):

@terenzreignz

terenzreignz (terenzreignz):

Well, two things... first \[\Large \sin(\alpha + \beta)= \sin(\alpha)\cos(\beta) + \cos(\alpha)\sin(\beta)\] and \[\Large \sin(3x) = \sin(2x+x)\]

OpenStudy (tkhunny):

\(\sin(3x) + \sin(x) = \sin(2x + x) + \sin(2x - x)\) might lead to something. You just have to play with it. You won't break it. Just try something.

OpenStudy (anonymous):

Ok @ terenzreignz

terenzreignz (terenzreignz):

Okay? So expand sin(3x) using the fact that sin(3x) = sin(2x + x) and sin(a+b) = sin(a)cos(b) + cos(a)sin(b) ^_^

OpenStudy (anonymous):

What are a and b?

OpenStudy (anonymous):

@terenzreignz

terenzreignz (terenzreignz):

Nothing, this is just a formula. If you want, you can have a = 2x and b = x and work from there...

OpenStudy (anonymous):

Ok, thank you so much.

terenzreignz (terenzreignz):

No problem :)